OpenStudy (anonymous):

0.25 r – 0.125 + 0.5 r = 0.5 + r solve for r please show steps so i can actually understand.

OpenStudy (anonymous):

add or subtract like terms 0.75r - 0.125 = 0.5 + r 0.25r= 0.625 r=2.5

OpenStudy (anonymous):

thanks